# Visual Hand Gesture Recognition with Deep Learning: A Comprehensive Review of Methods, Datasets, Challenges and Future Research Directions

### v2

@@ 5.2 Unimodal Pose-based Methods
Pose-based methods, also known as skeleton-based approaches, rely on structured motion data—such as skeleton graphs and pose heatmaps—to classify gestures. These methods often incorporate various forms of pose information, including joint coordinates, bones, motion trajectories, and joint angles, sometimes in combination (TMS-NET). They also explore joint subsets related to specific body regions (FDMSE-hierarchicalwindowedgraphattention), often adopting multi-stream architectures to handle these diverse inputs.
Given the natural graph structure of skeletal data, many methods employ Graph Convolutional Networks (GCNs) originally developed for human activity recognition. In the sections that follow, we group the approaches based on the structure of their backbone, following the same organization as in the previous subsection. Notably, pose information alone is often insufficient for accurate sign recognition (moryossef2021evaluating). Consequently, many models integrate additional cues—typically from RGB inputs—using 3DCNNs and perform score-level fusion to enhance performance.
@@ 5.2.1 Spatiotemporal
In SLU, pose-based methods model dynamic hand and body movements for gesture recognition. SKIM (SKIM) combines dynamic graph convolutional (GCNs) and temporal convolutional (TCNs) networks to analyze global and local motion, while MC-LSTM (ozdemir2023multi_MC-LSTM) fuses multiple streams (face, hand, body) via temporal convolutions and GCNs. P3D (P3D_ICCV) employs transformer-based part-wise and whole-body encoders, and HWGAT (FDMSE-hierarchicalwindowedgraphattention) enhances key body-part focus with hierarchical attention. TMS-Net (TMS-NET) introduces a six-stream GCN for skeleton, bone, and motion data with inter-stream fusion. To tackle data limitations, SLGTformer (SLGTFORMER) and VSCT (spoter_vsct) use transformers and few-shot learning, respectively. Dual-stream approaches such as (Jungpil-Shin) and MULTISTAGE-CGAR (MULTISTAGE-CGAR) integrate joint and motion data through temporal convolutions and attention. SwC GR-MMixer (SwC-GR-MMixer) combines GCNs with GRUs for temporal modeling and regularization. Other notable advances include L-TSSI-DenseNet (L-TSSI-DenseNet), which converts joint coordinates into image-like matrices for DenseNet processing; STST (STST), a two-stage transformer for local–global spatiotemporal features; and DSTA-SLR (DSTA_SLR_hu2024dynamic), which employs dynamic adjacency matrices and multi-stream fusion for improved recognition accuracy.
In HCI, pose-based methods address challenges such as data overfitting and the need for self-supervised learning frameworks. MAE-based frameworks (omar_ikne) use a masked autoencoder to reconstruct skeleton graphs, focusing on motion and spatial relationships. 3sISTGCN (JOINTS_MOTION_VECTORS_Improved_ST-GCN) proposes a three-stream model to process joints, bones, and motions, applying modified ST-GCN layers and late fusion for improved classification. Another approach, DG-STA (BOAT-MI_VR), tackles class-incremental learning in HCI by training models on non-overlapping tasks without access to old training data. DSTSA-GCN (DSTSA-CGN_cui2025dstsa) addresses temporal expression limitations by stacking three stages of graph convolutions and multiscale temporal convolutions. Finally, authors in (ST-CGNet-hu2025st) proposed ST-CGNet, a dual-stream architecture comprising a 3DCNN and a convolutional LSTM module to capture different levels of information. Attention mechanisms and an advanced fusion strategy are utilized to efficiently incorporate the features of each stream.
@@ 5.2.2 Spatial + Temporal
Many models use a two-step approach, where the spatial domain is first processed by a spatial perception module to extract spatial features. These features are then passed into a sequential model for temporal aggregation, resulting in a combined spatial-temporal representation of the input sequence. For instance, GCN-BERT (GCN-BERT) utilizes a two-layer GCN for spatial modeling, followed by BERT for temporal modeling.
Other general-purpose methods adopt a multi-stream, two-stage pose-based approach. In (SBI-DHGR), three streams, each containing three 1D convolutional layers, are combined with a parallel module of average pooling layers. The extracted features are then concatenated and input into an LSTM, followed by a fully connected layer and softmax for classification. Similarly, STr-GCN (STr-GCN) employs a two-stage model where a GCN spatial feature extractor is first applied to obtain intra-frame features and then passes the features through a transformer encoder for modeling inter-frame temporal relationships. The final classification scores are generated by an MLP.
@@ 5.2.3 Short-Term Spatiotemporal + Long-Term Temporal
For example, gesture recognition in autonomous driving has become an important research area. In (MG-GCT), the authors propose a pose-based approach called MG-GCT for recognizing police traffic gestures. The model uses two graph convolution streams: one to process skeleton joints and another to analyze joint motions. The features extracted from both streams are fused and passed through a transformer encoder for temporal modeling, with the final classification performed by an MLP. In another approach, TD-GCN (TD-GCN) introduces a pose-based method for IDHGR. Unlike traditional methods that use fixed skeleton topologies, TD-GCN employs temporal-sensitive topology learning to enhance feature quality. First, graph convolution is used to extract high-level spatiotemporal representations, which are then used to calculate temporal and channel-independent adjacency matrices. These matrices help model the skeleton, and the features are fused using the proposed TCF module. Authors in (ISLR-STGCN-attention-LSTM-ozdemir2025cross) proposed a dual-stream backbone processing both hand and whole body keypoints using ST-GCN models. Cross-attention is then applied to focus on the most important features, which are subsequently pass through a LSTM to model long-term dependencies.
@@ 5.2.4 Temporal
In (TELEOPERATED_SURGICAL_ROBOT), the authors proposed a pose-based method for surgical robot teleoperation. Preprocessed skeleton joints are fed into an LSTM, and the resulting features are passed to an MLP for final classification. The model was evaluated on a custom dataset. Similarly, in (Sabater_alberto), a cross-domain gesture recognition approach was developed using geometric features (e.g., bone angles), which are processed by a TCN. The resulting features are summarized into a sequence descriptor, and a linear classifier predicts the final class.
To address data scarcity, the authors in (ADVERSARIAL_DATA_AUGM) introduced AVSN, an adversarial learning-based data augmentation framework. The method jointly optimizes augmentation parameters and the HGR classifier. A BiLSTM serves as the discriminator, while an MLP acts as the generator. The generator produces challenging augmented data that exploits weaknesses in the discriminator, which in turn learns to correctly classify gestures.
@@ 5.2.5 Other
Several works explore alternative architectures or skeleton representations beyond the main categories. In (FS-EFCNN), a Fourier CNN with BiLSTM extracts spatiotemporal features from 3D joints for ISLR, using a trainable binary mask for feature selection. Similarly, (SCT-DCNN) applies feature thresholding based on variance and importance, followed by deep CNN classification. To overcome color-coded map limitations, (MRCNNSA) introduces the Joint Motion Affinity Map (JMAM) with a multi-stream attention-based 2DCNN for feature extraction and recognition.
Key frame selection is addressed in (KEYFRAMES_DRL_SKELETON) via FS-Net, a reinforcement learning agent that selects informative frames for DD-Net classification. Dual-stream architectures such as STGCN-LSTM (STGCN-LSTM-10965662) combine ST-GCN and self-attention ConvLSTM for capturing global and local features, while (Yangke_patcog) proposes a lightweight variant merging attention-based and autoencoder streams. Finally, the Hierarchical Self-Attention Network (HAN) (HAN-liu2025han) employs modular attention across finger joints (J-Att, F-Att, T-Att, Fusion-Att) for efficient spatial–temporal modeling, achieving strong accuracy with just 0.05 GFLOPs.
@@ 5.3.1 General Purpose
In (Yunan_Li), a dual-stream 3DCNN with average fusion was proposed for RGBD gesture recognition, using a memory bank to retain discriminative features ($z$) while filtering redundant ones ($\upsilon$) by maximizing mutual information with class labels. Similarly, (RAAR3DNet) employed Neural Architecture Search (NAS) with Dynamic Static Attention (DSA) modules and a modified I3D backbone to emphasize motion and hand regions, supported by an auxiliary HeatmapNet and late RGBD fusion.
To better exploit modality interactions, (Benjia_Zhou) introduced a decoupling–recoupling framework combining independent spatial and temporal modeling with mid-level fusion through a shared MLP-AE module. Multi-level fusion was further explored in (CTFB_Hampiholi), integrating 3DCNNs with a Convolutional Transformer Fusion Block (CTFB) that fuses intermediate cross-stream features.
To address data scarcity, several works explored zero-shot learning (ZSL) by incorporating textual gesture descriptions encoded by pretrained language models (mainly BERT). In (ZS-GR), a dual-stream model combined a ViT for spatial modeling and an LSTM-MLP for temporal modeling, learning visual features aligned with textual embeddings. Similarly, (ZS-DHGR) fused skeleton-derived geometric features with RGB-C3D-LSTM-AE representations into a shared semantic space for classification. To reduce computational cost, (ConvMixFormer) replaced the transformer’s self-attention mechanism with convolutional layers and a gated feedforward network, resulting in the lightweight ConvMixFormer architecture.
Knowledge distillation and contrastive learning were applied in (3STREAM_DISTILLATION), where three 3DCNNs (RGB, Depth, Optical Flow) were fused with Coordinate Attention, using I3D as teacher and ResNet as student, while integrating CLIP for zero-shot recognition. Finally, (MFST) and (3D_CDC_NAS_TIP_2021) replaced standard convolutions with 3D Central Difference Convolution (3D-CDC), combining hybrid residual blocks, transformers, and NAS-based optimization for enhanced spatiotemporal modeling.
@@ 5.3.2 Sign Language Understanding
Recent works on Isolated Sign Language Recognition (ISLR) have focused on multimodal, pose-centric architectures combining RGB and skeleton data. SignBERT (SignBERT_ICCV) employs a spatial GCN with temporal and hand chirality encoding, feeding features into a transformer encoder and a hand-model-aware decoder for self-supervised pretraining, with parallel RGB predictions fused at inference. Its enhanced variant, SignBERT+ (SignBERT+), adds spatial-temporal position encoding and an improved RGB–pose fusion module (Fig. 6). Similarly, MASA (MASA) uses the same GCN+Transformer backbone with contrastive learning on pose reconstruction, while BEST (BEST) embeds pose triplets with BERT and applies late fusion with RGB data.
Figure 6: The pose-based stream of SignBERT+ (SignBERT+), an extension of SignBERT (SignBERT_ICCV), first uses a GCN to extract gesture features, which are combined with spatial-temporal position encoding and processed by a transformer encoder. During self-supervised pretraining, a MANO-based decoder reconstructs masked inputs, and is replaced by a prediction head for downstream fine-tuning.
To improve feature invariance, (Weichao-Zhao) applies a contrastive approach enforcing consistency among hand, trunk, and motion features, using the same GCN+Transformer backbone with late fusion of pose and RGB predictions. SAM-SLR (SAM_SLR) fuses six streams (RGB, depth, skeleton, optical flow, pose heatmaps, HHA), while SAM-SLR-v2 (jiang2021sign-sam-slr-v2) adds a 3D hand graph stream and learnable fusion weights.
VTN-HC/PF (VTN-HC-PF) processes hand-cropped images via a 2DCNN, transformer, and MLP for final predictions, with PF incorporating pose-based geometrical features. Motion History Images are used in (MOTION_HISTORY_Mercanoglu), analyzed by a 3DCNN with attention, and in a dual-stream setup combined with 2DCNN RGB features via weighted summation.
To tackle data scarcity, ZSSLR (ZSSLR) uses a CLIP-based zero-shot framework, while (FSSLR) proposes a few-shot cross-lingual method combining two 3DCNNs and a pose estimator with a Temporal Feature Aggregation Module. NLA-SLR (NLA_SLR) targets semantically similar signs with a multistream S3D backbone and language-aware label smoothing. MSNN (MSNN) uses five I3D streams plus ST-GCN, and hDNN-SLR (hDNN-SLR) combines a 3DCNN, attention-based BiLSTM, and a VAE-based module. Finally, Uni-Sign (UNISIGNli2025) proposes a unified multimodal architecture with three pose encoders and an RGB encoder fused via attention, followed by temporal encoding and LLM-based pretraining for ISLR and CSLR.
@@ 5.3.3 Human-Computer Interaction
Motivated by the difficulty of recognizing fine-grained gestures in HCI applications, authors in (HandSense) proposed HandSense, a method relying on depth and RGB data. Two 3DCNNs are deployed to extract features for each one modality. These features are fused and fed into a SVM to obtain the final prediction. Authors in (mXception_and_RDDI) proposed an approach relying on RGBD Dynamic Images (RDDIs). These images, are obtained by the combination of RGB and estimated depth data. The resulted RDDI is fed into a modified Xception, resulting from removing the top 4 Xception modules, in order to extract meaningful features. These features are finally used to perform the classification.
@@ 5.3.4 Other
Inspired by the success of transformers, GestFormer (GestFormer) introduces a lightweight transformer-based framework for gesture-driven HRI. A 2DCNN extracts per-frame features, processed by PoolFormer layers (replacing attention) and two custom blocks for multiscale pooling/WT and token mixing. Each modality is processed independently, and outputs are averaged. In (DHGR_3D_POSE_HRI), a hybrid 3DCNN/LSTM model fuses RGB, depth, and skeleton data (from an enhanced pose estimator) via weighted sum. The fused input feeds a spatiotemporal model combining 2D/3D convolutions and ConvLSTM layers. Authors in (MDSI-fang2025mdsi) emphasize on two major issues, information redundancy and the difficulty to recognize similar gestures. Their approach leverages a CLIP encoder to provide qualitative embeddings and enhance the feature extractor through label smoothing as well as a semantic filter to incorporate multimodal data.

### v3

@@ 5.2 Unimodal Pose-based Methods
Pose-based methods, also known as skeleton-based approaches, rely on structured motion data—such as skeleton graphs and pose heatmaps—to classify gestures. These methods often incorporate various forms of pose information, including joint coordinates, bones, motion trajectories, and joint angles, sometimes in combination (40). They also explore joint subsets related to specific body regions (152), often adopting multi-stream architectures to handle these diverse inputs.
Given the natural graph structure of skeletal data, many methods employ Graph Convolutional Networks (GCNs) originally developed for human activity recognition. In the sections that follow, we group the approaches based on the structure of their backbone, following the same organization as in the previous subsection. Notably, pose information alone is often insufficient for accurate sign recognition (139). Consequently, many models integrate additional cues—typically from RGB inputs—using 3DCNNs and perform score-level fusion to enhance performance.
@@ 5.2.1 Spatiotemporal
In SLU, pose-based methods model dynamic hand and body movements for gesture recognition. SKIM (115) combines dynamic graph convolutional (GCNs) and temporal convolutional (TCNs) networks to analyze global and local motion, while MC-LSTM (148) fuses multiple streams (face, hand, body) via temporal convolutions and GCNs. P3D (109) employs transformer-based part-wise and whole-body encoders, and HWGAT (152) enhances key body-part focus with hierarchical attention. TMS-Net (40) introduces a six-stream GCN for skeleton, bone, and motion data with inter-stream fusion. To tackle data limitations, SLGTformer (187) and VSCT (26) use transformers and few-shot learning, respectively. Dual-stream approaches such as (180) and MULTISTAGE-CGAR (134) integrate joint and motion data through temporal convolutions and attention. SwC GR-MMixer (122) combines GCNs with GRUs for temporal modeling and regularization. Other notable advances include L-TSSI-DenseNet (108), which converts joint coordinates into image-like matrices for DenseNet processing; STST (222), a two-stage transformer for local–global spatiotemporal features; and DSTA-SLR (86), which employs dynamic adjacency matrices and multi-stream fusion for improved recognition accuracy.
In HCI, pose-based methods address challenges such as data overfitting and the need for self-supervised learning frameworks. MAE-based frameworks (88) use a masked autoencoder to reconstruct skeleton graphs, focusing on motion and spatial relationships. 3sISTGCN (186) proposes a three-stream model to process joints, bones, and motions, applying modified ST-GCN layers and late fusion for improved classification. Another approach, DG-STA (5), tackles class-incremental learning in HCI by training models on non-overlapping tasks without access to old training data. DSTSA-GCN (32) addresses temporal expression limitations by stacking three stages of graph convolutions and multiscale temporal convolutions. Finally, authors in (81) proposed ST-CGNet, a dual-stream architecture comprising a 3DCNN and a convolutional LSTM module to capture different levels of information. Attention mechanisms and an advanced fusion strategy are utilized to efficiently incorporate the features of each stream.
@@ 5.2.2 Spatial + Temporal
Many models use a two-step approach, where the spatial domain is first processed by a spatial perception module to extract spatial features. These features are then passed into a sequential model for temporal aggregation, resulting in a combined spatial-temporal representation of the input sequence. For instance, GCN-BERT (194) utilizes a two-layer GCN for spatial modeling, followed by BERT for temporal modeling.
Other general-purpose methods adopt a multi-stream, two-stage pose-based approach. In (143), three streams, each containing three 1D convolutional layers, are combined with a parallel module of average pooling layers. The extracted features are then concatenated and input into an LSTM, followed by a fully connected layer and softmax for classification. Similarly, STr-GCN (184) employs a two-stage model where a GCN spatial feature extractor is first applied to obtain intra-frame features and then passes the features through a transformer encoder for modeling inter-frame temporal relationships. The final classification scores are generated by an MLP.
@@ 5.2.3 Short-Term Spatiotemporal + Long-Term Temporal
For example, gesture recognition in autonomous driving has become an important research area. In (65), the authors propose a pose-based approach called MG-GCT for recognizing police traffic gestures. The model uses two graph convolution streams: one to process skeleton joints and another to analyze joint motions. The features extracted from both streams are fused and passed through a transformer encoder for temporal modeling, with the final classification performed by an MLP. In another approach, TD-GCN (120) introduces a pose-based method for IDHGR. Unlike traditional methods that use fixed skeleton topologies, TD-GCN employs temporal-sensitive topology learning to enhance feature quality. First, graph convolution is used to extract high-level spatiotemporal representations, which are then used to calculate temporal and channel-independent adjacency matrices. These matrices help model the skeleton, and the features are fused using the proposed TCF module. Authors in (149) proposed a dual-stream backbone processing both hand and whole body keypoints using ST-GCN models. Cross-attention is then applied to focus on the most important features, which are subsequently pass through a LSTM to model long-term dependencies.
@@ 5.2.4 Temporal
In (156), the authors proposed a pose-based method for surgical robot teleoperation. Preprocessed skeleton joints are fed into an LSTM, and the resulting features are passed to an MLP for final classification. The model was evaluated on a custom dataset. Similarly, in (170), a cross-domain gesture recognition approach was developed using geometric features (e.g., bone angles), which are processed by a TCN. The resulting features are summarized into a sequence descriptor, and a linear classifier predicts the final class.
To address data scarcity, the authors in (142) introduced AVSN, an adversarial learning-based data augmentation framework. The method jointly optimizes augmentation parameters and the HGR classifier. A BiLSTM serves as the discriminator, while an MLP acts as the generator. The generator produces challenging augmented data that exploits weaknesses in the discriminator, which in turn learns to correctly classify gestures.
@@ 5.2.5 Other
Several works explore alternative architectures or skeleton representations beyond the main categories. In (1), a Fourier CNN with BiLSTM extracts spatiotemporal features from 3D joints for ISLR, using a trainable binary mask for feature selection. Similarly, (10) applies feature thresholding based on variance and importance, followed by deep CNN classification. To overcome color-coded map limitations, (100) introduces the Joint Motion Affinity Map (JMAM) with a multi-stream attention-based 2DCNN for feature extraction and recognition.
Key frame selection is addressed in (51) via FS-Net, a reinforcement learning agent that selects informative frames for DD-Net classification. Dual-stream architectures such as STGCN-LSTM (68) combine ST-GCN and self-attention ConvLSTM for capturing global and local features, while (111) proposes a lightweight variant merging attention-based and autoencoder streams. Finally, the Hierarchical Self-Attention Network (HAN) (119) employs modular attention across finger joints (J-Att, F-Att, T-Att, Fusion-Att) for efficient spatial–temporal modeling, achieving strong accuracy with just 0.05 GFLOPs.
@@ 5.3.1 General Purpose
In (112), a dual-stream 3DCNN with average fusion was proposed for RGBD gesture recognition, using a memory bank to retain discriminative features ($z$) while filtering redundant ones ($\upsilon$) by maximizing mutual information with class labels. Similarly, (227) employed Neural Architecture Search (NAS) with Dynamic Static Attention (DSA) modules and a modified I3D backbone to emphasize motion and hand regions, supported by an auxiliary HeatmapNet and late RGBD fusion.
To better exploit modality interactions, (228) introduced a decoupling–recoupling framework combining independent spatial and temporal modeling with mid-level fusion through a shared MLP-AE module. Multi-level fusion was further explored in (66), integrating 3DCNNs with a Convolutional Transformer Fusion Block (CTFB) that fuses intermediate cross-stream features.
To address data scarcity, several works explored zero-shot learning (ZSL) by incorporating textual gesture descriptions encoded by pretrained language models (mainly BERT). In (164), a dual-stream model combined a ViT for spatial modeling and an LSTM-MLP for temporal modeling, learning visual features aligned with textual embeddings. Similarly, (162) fused skeleton-derived geometric features with RGB-C3D-LSTM-AE representations into a shared semantic space for classification. To reduce computational cost, (58) replaced the transformer’s self-attention mechanism with convolutional layers and a gated feedforward network, resulting in the lightweight ConvMixFormer architecture.
Knowledge distillation and contrastive learning were applied in (198), where three 3DCNNs (RGB, Depth, Optical Flow) were fused with Coordinate Attention, using I3D as teacher and ResNet as student, while integrating CLIP for zero-shot recognition. Finally, (127) and (215) replaced standard convolutions with 3D Central Difference Convolution (3D-CDC), combining hybrid residual blocks, transformers, and NAS-based optimization for enhanced spatiotemporal modeling.
@@ 5.3.2 Sign Language Understanding
Recent works on Isolated Sign Language Recognition (ISLR) have focused on multimodal, pose-centric architectures combining RGB and skeleton data. SignBERT (78) employs a spatial GCN with temporal and hand chirality encoding, feeding features into a transformer encoder and a hand-model-aware decoder for self-supervised pretraining, with parallel RGB predictions fused at inference. Its enhanced variant, SignBERT+ (77), adds spatial-temporal position encoding and an improved RGB–pose fusion module (Fig. 6). Similarly, MASA (223) uses the same GCN+Transformer backbone with contrastive learning on pose reconstruction, while BEST (224) embeds pose triplets with BERT and applies late fusion with RGB data.
Figure 6: The pose-based stream of SignBERT+ (77), an extension of SignBERT (78), first uses a GCN to extract gesture features, which are combined with spatial-temporal position encoding and processed by a transformer encoder. During self-supervised pretraining, a MANO-based decoder reconstructs masked inputs, and is replaced by a prediction head for downstream fine-tuning.
To improve feature invariance, (225) applies a contrastive approach enforcing consistency among hand, trunk, and motion features, using the same GCN+Transformer backbone with late fusion of pose and RGB predictions. SAM-SLR (91) fuses six streams (RGB, depth, skeleton, optical flow, pose heatmaps, HHA), while SAM-SLR-v2 (90) adds a 3D hand graph stream and learnable fusion weights.
VTN-HC/PF (38) processes hand-cropped images via a 2DCNN, transformer, and MLP for final predictions, with PF incorporating pose-based geometrical features. Motion History Images are used in (133), analyzed by a 3DCNN with attention, and in a dual-stream setup combined with 2DCNN RGB features via weighted summation.
To tackle data scarcity, ZSSLR (147) uses a CLIP-based zero-shot framework, while (25) proposes a few-shot cross-lingual method combining two 3DCNNs and a pose estimator with a Temporal Feature Aggregation Module. NLA-SLR (235) targets semantically similar signs with a multistream S3D backbone and language-aware label smoothing. MSNN (130) uses five I3D streams plus ST-GCN, and hDNN-SLR (159) combines a 3DCNN, attention-based BiLSTM, and a VAE-based module. Finally, Uni-Sign (113) proposes a unified multimodal architecture with three pose encoders and an RGB encoder fused via attention, followed by temporal encoding and LLM-based pretraining for ISLR and CSLR.
@@ 5.3.3 Human-Computer Interaction
Motivated by the difficulty of recognizing fine-grained gestures in HCI applications, authors in (221) proposed HandSense, a method relying on depth and RGB data. Two 3DCNNs are deployed to extract features for each one modality. These features are fused and fed into a SVM to obtain the final prediction. Authors in (96) proposed an approach relying on RGBD Dynamic Images (RDDIs). These images, are obtained by the combination of RGB and estimated depth data. The resulted RDDI is fed into a modified Xception, resulting from removing the top 4 Xception modules, in order to extract meaningful features. These features are finally used to perform the classification.
@@ 5.3.4 Other
Inspired by the success of transformers, GestFormer (56) introduces a lightweight transformer-based framework for gesture-driven HRI. A 2DCNN extracts per-frame features, processed by PoolFormer layers (replacing attention) and two custom blocks for multiscale pooling/WT and token mixing. Each modality is processed independently, and outputs are averaged. In (54), a hybrid 3DCNN/LSTM model fuses RGB, depth, and skeleton data (from an enhanced pose estimator) via weighted sum. The fused input feeds a spatiotemporal model combining 2D/3D convolutions and ConvLSTM layers. Authors in (48) emphasize on two major issues, information redundancy and the difficulty to recognize similar gestures. Their approach leverages a CLIP encoder to provide qualitative embeddings and enhance the feature extractor through label smoothing as well as a semantic filter to incorporate multimodal data.